# Simple and Provable Scaling Laws for the Test-Time Compute of Large Language Models

### v4

@@ 1 Introduction
Despite the astonishing advancements of large language models (LLMs) in the past few years, they still face challenges with reliability and stability. This hinders their applications in high-stakes scenarios where a problem need to be solved with success probability $99.9\%$ rather than $90\%$. Similarly, in an LLM-based agentic workflow that involves solving many sub-problems, each of them need to be solved with near-perfect success rate, since a single error in the process can lead to an incorrect final output. In these and many other similar scenarios, one is willing to boost the success probability by spending more test-time compute on LLM inference. One category of methods for this purpose include iterative approaches like generating a sequential chain of thoughts [44, 18, 29, 53, 7, 38] or self-refinement [6, 25, 8, 50, 51]. Another category, which is the focus of this work, is about repeatedly sampling multiple solutions and then aggregating them for the final output [3, 42, 2, 21, 13, 15, 37, 5, 22, 55, 48, 54, 35]. These two categories are complementary and often used together for the best performance [36, 45, 53, 12, 28].
@@
Strong baseline methods widely adopted in practice may fail (in theory and practice) to achieve this goal, even if a single LLM call already solves the input problem correctly with high or moderate success probability. For example, best-of-N (BoN) sampling with an imperfect verifier might suffer from performance decay as the number of sampled solutions grows because, as explained in Section 5.1 of [5], “the benefits of search are eventually outweighed by the risk of finding adversarial solutions that fool the verifier”. Indeed, prior works highlighted that developing test-time scaling methods that do not rely on perfect verifiers remains an important direction for further research [2, 37]. Majority voting, another strong baseline, might fail for different reasons: even if one LLM call has moderate probability, say 45%, of generating a correct final answer, the success probability of majority voting will actually converge to zero as the number of samples grow if there exists an incorrect final answer that a single LLM call generates with probability 46% [3].
@@
In pursuit of provable inference scaling laws, we propose a two-stage knockout-style algorithm that first generates multiple candidate solutions, and then select one via a knockout tournament where pairwise comparisons among the candidates are conducted. We prove theoretically that its failure probability decays to zero exponentially (Theorem 1) or by a power law (Theorem 2) with respect to the total number of LLM calls, depending on the specific way of scaling. These guarantees rely on two natural (and arguably minimal) assumptions: (1) the LLM can generate a correct solution with non-zero probability, and (2) the LLM can do better than a random guess in choosing the right winner between a pair of correct and incorrect solutions.
We further propose a two-stage league-style algorithm that also enjoys a provable scaling law. Unlike the knockout-style algorithm that eliminates a candidate upon loss to a single opponent, the league-style algorithm evaluates each candidate by its average win rate against multiple opponents. We prove that its failure probability also decays to zero exponentially (Theorem 3), under technical assumptions that are analogous to but more robust than those required by the knockout-style algorithm.
Both proposed algorithms require a black-box LLM and nothing else (e.g., no external verifier or reward model) for a minimalistic implementation, which makes them appealing for practical applications and easy to adapt for different scenarios. Our practical implementations are efficient and scalable, with support for parallel and distributed computation. Empirical results with diverse LLMs (Llama3.1, Qwen2.5, GPT-4o, QwQ-32B) and datasets (GPQA, MMLU-Pro, MATH-500) validate our theories and demonstrate the outstanding scaling properties of both algorithms.
@@ 2 A two-stage knockout-style algorithm
    Generation. We first generate $N$ candidate solutions, which can run in parallel. In situations where the final answer contains only a few tokens (e.g., for multiple-choice problems or math calculation), we require that each solution contains a thinking process, which can be elicited by chain-of-thought (CoT) prompting [44, 18] for example; such information can be useful for enhancing pairwise comparisons in the next stage.
@@
Figure 1 visualizes this process. For a minimalistic implementation, both stages can be executed with a single black-box LLM (or an ensemble of multiple LLMs). We next introduce some formal notations, followed by our analysis for both success probability and computational efficiency [16, 4] of the proposed algorithm, to be presented in Sections 2.1 and 2.2 respectively.
@@ 2.1 Analysis of success probability
Theorem 1.
If Assumption 1 holds for the input problem, then the probability that the proposed knockout-style algorithm returns an incorrect final output decays to zero exponentially with respect to the hyperparameters $N$ and $K$:
@@
A proof can be found in Appendix A.1. Another way to interpret this theorem is as follows: for a targeted success probability $1-\delta$ (which can be arbitrarily close to 1 as the failure probability $\delta>0$ approaches zero), it suffices to have
@@
Although Theorem 1 guarantees an arbitrarily small failure probability, it requires $K$ to be sufficiently large, depending on the value of $p_{\text{comp}}$ that might be unknown a priori in practice. To resolve this, we provide an alternative theorem suggesting that scaling up $N$ alone is sufficient, even when $K$ is a fixed constant and thus there is still a good chance that the wrong winner is identified when comparing a specific pair of candidates.
To streamline the statement of our theorem, we introduce the notations $\mathcal{M}_{\text{comp},K}$ and $p_{\text{comp},K}$, which generalize $\mathcal{M}_{\text{comp}}$ and $p_{\text{comp}}$ that appear in Assumption 1. Let $\mathcal{M}_{\text{comp},K}$ denote the probability distribution of the comparison result obtained with $K$ independent LLM calls followed by majority voting (with ties broken randomly). Then we have
@@
Theorem 2.
Suppose that Assumption 1 holds and $N$ is a power of 2. Let $p_{i}$ be the probability that a candidate solution at the $i$-th level of the knockout tournament is correct, where $i=0,1,\dots,\log_{2}N$111 In our notations, the zeroth level of the tournament contains $N$ initial candidates, the first level contains $N/2$ winners after the first round of pairwise comparisons, and so on. All candidates within the same level of the knockout tournament have the same probability of being a correct one, due to their symmetric roles. Then $p_{0}=p_{\text{gen}}$ and
@@
Consequently, the success probability of the overall algorithm, namely $p_{\log_{2}N}$, converges to 1 as $N$ grows; for any $0.50<\delta<$ 0.5, one has $p_{\log_{2}N}\geq 1-\delta$ as long as
@@
A proof can be found in Appendix A.2. The linear relationship between $\log_{2}N$ and $\log(1/\delta)$ reveals a power-law relationship between the failure probability $\delta$ and the number of candidates $N$.
@@ 3 A two-stage league-style algorithm
Recall that Assumption 1 in the previous section requires $p_{\text{comp}}>0.5$ for any pair of correct and incorrect solutions, which renders it somewhat restricted and non-robust. This motivates us to propose a two-stage league-style algorithm that also enjoys a provable inference scaling law, under technical assumptions analogous to but more robust than those required by the knockout-style algorithm.
@@
To begin with, we generate $N$ candidate solutions $y_{1},\dots,y_{N}\sim\mathcal{M}_{\text{gen}}(x)$ as before. Then, for each candidate with index $i\in[N]$, we randomly sample $K$ opponents with indices $o_{i}(1),\dots,o_{i}(K)\in[N]\backslash\{i\}$ uniformly and with replacement, conduct one independent pairwise comparison against each opponent, and obtain the responses $r_{i}(j)\sim\mathcal{M}_{\text{comp}}(x,y_{i},y_{o_{i}(j)}),j\in[K]$. The average win rate of each candidate $y_{i}$ is estimated by
$\widehat{\mu}_{i}\coloneqq\frac{1}{K}\sum_{j\in[K]}\phi\big{(}r_{i}(j),y_{i},y_{o_{i}(j)}\big{)},$
where $\phi(r_{i}(j),y_{i},y_{o_{i}(j)})$ denotes the score assigned, based on $r_{i}(j)$, to the candidate $y_{i}$ in its comparison against $y_{o_{i}(j)}$, e.g., 1 for a win, 0 for a loss, and 0.5 for a tie. Finally, the candidate with the highest average win rate $\widehat{\mu}_{i}$ is chosen (with ties broken randomly) as the output of the algorithm.
@@
In other words, we assume that the LLM can generate a correct-and-strong solution with non-zero probability, whose average win rate against $\mathcal{M}_{\text{gen}}$ is higher than that of any incorrect solution. Note that Assumption 2 can be tolerant of systematic errors by LLMs in comparing certain pairs of candidates, i.e., it may still hold true when there exist a correct solution $y$ and incorrect solution $y^{\prime}$ such that ${\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime})}\phi(r,y,y^{\prime})<0.5$, whereas Assumption 1 fails in such cases.
Intuitively, if Assumption 2 holds true and the hyperparameters $N$ and $K$ are sufficiently large, then with high probability, (1) there exist initial candidates that are correct-and-strong solutions, and (2) $\widehat{\mu}_{i}$ is an accurate estimate of $\mu_{y_{i}}$ for each $i\in[N]$. These conditions together lead to a correct final output of the algorithm. We formalize this intuition in the following theorem.
Theorem 3.
If Assumption 2 holds for the input problem, then the probability that the league-style algorithm (with hyperparameters $N$ and $K$) returns an incorrect final output is bounded by
@@
This theorem, whose proof can be found in Appendix A.3, ensures that the failure probability of the league-style algorithm decays to zero exponentially with respect to $N$ and $K$. Another way to interpret this theorem is as follows: to guarantee success probability $1-\delta$, it suffices to have
@@
Remark 1.
The provable success of the knockout-style algorithm relies on Assumption 1, while that of the league-style algorithm relies on Assumption 2. Although the latter is conceptually more robust than the former, we note that neither assumption is strictly weaker than the other (and thus both algorithms have their unique values). In other words, there exist scenarios where Assumption 1 holds true while Assumption 2 does not, and also scenarios where the reverse is true. Interested readers may refer to Appendix B for some minimal examples.
@@ 4 Experiments
We use three datasets for our experiments: GPQA [34], MMLU-Pro [43] and MATH-500 [26]. GPQA consists of over 1000 graduate-level multiple-choice questions splitted into three categories (“main”, “diamond” and “extended”), all of which are used in our experiments. MMLU-Pro contains 14 categories of multiple-choice questions, some of which require advanced reasoning while others are more knowledge-focused. Due to limited computational resources, we use a randomly sampled subset of 100 questions for each category of MMLU-Pro in our experiments, which leads to a total of 1400 questions; we refer to this subset as MMLU-Pro-S throughout this work. MATH-500 is a subset of 500 problems from the MATH dataset introduced in [22]. We focus mainly on GPQA in this section, deferring empirical results for MMLU-Pro-S and MATH-500 (as well as supplementary results for GPQA) to Appendix C.
@@
We use Llama3.1-70B-Instruct (Llama3.1 for short) [24] and Qwen2.5-72B-Instruct (Qwen2.5 for short) [31] in our experiments, as well as a Mixed option that uses a mixture of both LLMs [40, 54, 13]: during the generation stage, half of the initial candidates are sampled by Llama3.1 and the other half by Qwen2.5; similarly, when a pair of candidates are compared for multiple times during the aggregation stage, half of them are done by Llama3.1 and the other half by Qwen2.5. The rationale is that the capabilities of different LLMs can be complementary to some extent, and thus using a mixture of them can make it more likely that Assumptions 1 and 2 hold true222 To formalize this intuition, consider a minimal scenario with two LLMs denoted by $M_{1}$ and $M_{2}$, and two problems denoted by $x_{1}$ and $x_{2}$. Suppose that $M_{1}$ is effective for the first problem $x_{1}$ (with $p_{\text{gen}}=0.2$ and $p_{\text{comp}}=0.7$) but ineffective for $x_{2}$ (with $p_{\text{gen}}=0$ and $p_{\text{comp}}=0.5$), while the reverse holds true for $M_{2}$. When either LLM is used alone, only one problem satisfies Assumption 1. However, when a mixture of two LLMs is used, both problems now satisfy Assumption 1 with $p_{\text{gen}}=(0+0.2)/2=0.1>0$ and $p_{\text{comp}}=(0.5+0.7)/2=0.6>0.5$, and thus can be solved by our algorithm. . Other models considered in our experiments include QwQ-32B [32], a long-CoT reasoning LLM, and GPT-4o [10], a proprietary API-based LLM; due to high computational or monetary costs, they are tested only on GPQA-diamond for a smaller range of $N$.
We leverage zero-shot chain-of-thought prompting [18] for both generation and aggregation stages of the proposed algorithms. Unless specified otherwise, for the knockout-style algorithm, we fix $K=4$ for Llama3.1/ Qwen2.5/ Mixed, and $K=2$ for GPT-4o/ QwQ-32B; for the league-style algorithm, we consider a round-robin [47] version of it, with $K=4$ comparisons conducted between each of (${N\choose 2}$ pairs of initial candidates. To make the proposed algorithms efficient and scalable in practice, we implement them based on AgentScope [9], a multi-agent framework that supports parallel and distributed computation. Further implementation details can be found in Appendix C.1.

### v5

# Provable Scaling Laws for the Test-Time Compute of Large Language Models

@@ 1 Introduction
Despite the astonishing advancements of large language models (LLMs) in the past few years, they still face challenges with reliability and stability. This hinders their applications in high-stakes scenarios where a problem need to be solved with success probability $99.9\%$ rather than $90\%$. Similarly, in an LLM-based agentic workflow that involves solving many sub-problems, each of them need to be solved with near-perfect success rate, since a single error in the process can lead to an incorrect final output. In these and many other similar scenarios, one is willing to boost the success probability by spending more test-time compute on LLM inference. One category of methods for this purpose include iterative approaches like generating a sequential chain of thoughts Wei2022ChainOT ; Kojima2022ZeroShot ; openai2024o1 ; zeng2024scalingsearchlearningroadmap ; deepseekai2025deepseekr1incentivizingreasoningcapability ; kimiteam2025kimik15scalingreinforcement or self-refinement davis2024networksnetworkscomplexityclass ; madaan2023selfrefine ; du2023improvingfactualityreasoninglanguage ; yinetal2023exchange ; yinetal2024aggregation . Another category, which is the focus of this work, is about repeatedly sampling multiple solutions and then aggregating them for the final output; examples include best-of-N sampling cobbe2021trainingverifierssolvemath ; brown2024largelanguagemonkeysscaling ; stroebl2024inferencescalingflawslimits ; zhang2024generative ; schaeffer2025largelanguagemonkeyspower , majority voting wang2023selfconsistency ; chen2024llm ; li2024more , among others jiang2023llmblender ; kambhampati2024modulo ; lightman2024lets ; wu2024empiricalanalysiscomputeoptimalinference ; zhang2024scalingllminferenceoptimized . These two categories are complementary and often used together for the best performance snell2024scalingllmtesttimecompute ; welleck2024from ; zeng2024scalingsearchlearningroadmap ; ji2025testtimecomputingsystem1thinking ; muennighoff2025s1simpletesttimescaling .
@@
Strong baseline methods widely adopted in practice may fail (in theory and practice) to achieve this goal, even if a single LLM call already solves the input problem correctly with high or moderate success probability. For example, best-of-N (BoN) sampling with an imperfect verifier might suffer from performance decay as the number of sampled solutions grows because, as explained in Section 5.1 of cobbe2021trainingverifierssolvemath , “the benefits of search are eventually outweighed by the risk of finding adversarial solutions that fool the verifier”. Indeed, prior works highlighted that developing test-time scaling methods that do not rely on perfect verifiers remains an important direction for further research brown2024largelanguagemonkeysscaling ; stroebl2024inferencescalingflawslimits . Majority voting, another strong baseline, might fail for different reasons: even if the LLM has moderate probability, say 45%, of generating a correct final answer, the success probability of majority voting will actually converge to zero as the number of samples grow if there exists an incorrect final answer that a single LLM call generates with probability 46% chen2024llm .
@@
In pursuit of provable inference scaling laws, we propose a two-stage knockout-style algorithm that first generates multiple candidate solutions, and then select one via a knockout tournament where pairwise comparisons among the candidates are conducted. We prove theoretically that its failure probability decays to zero exponentially (Theorem 2.3) or by a power law (Theorem 2.4) with respect to the total number of LLM calls, depending on the specific way of scaling. These guarantees rely on two assumptions: (1) the LLM can generate a correct solution with non-zero probability, and (2) the LLM can do better than a random guess in choosing the right winner between any pair of correct and incorrect solutions.
We further propose a two-stage league-style algorithm that also enjoys a provable scaling law. Unlike the knockout-style algorithm that eliminates a candidate upon loss to a single opponent, the league-style algorithm evaluates each candidate by its average win rate against multiple opponents. We prove that its failure probability also decays to zero exponentially (Theorem 3.3) as its test-time compute scales up, under the technical assumption that there exist correct solutions whose average win rates against a distribution of opponents are higher than that of any incorrect solution.
Both proposed algorithms require a black-box LLM and nothing else (e.g., no external verifier or reward model) for a minimalistic implementation, which makes them appealing for practical applications and easy to adapt for different scenarios. Our practical implementations are efficient and scalable, with support for parallel and distributed computation. While the technical assumptions in our theories might seem strong from a practical perspective, our empirical results confirm that the proposed algorithms — developed based on the theoretical insights — indeed perform well and demonstrate outstanding scaling properties across diverse LLMs (Llama3.1, Qwen2.5, GPT-4o, QwQ-32B) and datasets (GPQA, MMLU-Pro, MATH-500).
@@ 2 A two-stage knockout-style algorithm
    Generation. We first generate $N$ candidate solutions, which can run in parallel. In situations where the final answer contains only a few tokens (e.g., for multiple-choice problems or math calculation), we require that each solution contains a thinking process, which can be elicited by chain-of-thought (CoT) prompting Wei2022ChainOT ; Kojima2022ZeroShot for example; such information can be useful for enhancing pairwise comparisons in the next stage.
@@
Figure 1 visualizes this process. For a minimalistic implementation, both stages can be executed with a single black-box LLM (or an ensemble of multiple LLMs). We next introduce some formal notations, followed by our analysis for both success probability and computational efficiency kapoor2024aiagentsmatter ; chen2025designing of the proposed algorithm, to be presented in Sections 2.1 and 2.2 respectively.
@@ 2.1 Analysis of success probability
Remark 2.2.
While this assumption seems minimal and natural at first glance, its requirement that $p_{\text{comp}}>0.5$ holds for any pair of correct and incorrect solutions renders it somewhat restricted and non-robust. This will motivate our development for an alternative algorithm and its provable scaling law under different technical assumptions, to be elaborated in Section 3.
@@
Theorem 2.3.
If Assumption 2.1 holds for the input problem, then the probability that the proposed knockout-style algorithm returns an incorrect final output decays to zero exponentially with respect to the hyperparameters $N$ and $K$:
@@
A proof can be found in Appendix B.1. Another way to interpret this theorem is as follows: for a targeted success probability $1-\delta$ (which can be arbitrarily close to 1 as the failure probability $\delta>0$ approaches zero), it suffices to have
@@
Although Theorem 2.3 guarantees an arbitrarily small failure probability, it requires $K$ to be sufficiently large, depending on the value of $p_{\text{comp}}$ that might be unknown a priori in practice. To resolve this, we provide an alternative theorem suggesting that scaling up $N$ alone is sufficient, even when $K$ is a fixed constant and thus there is still a good chance that the wrong winner is identified when comparing a specific pair of candidates.
To streamline the statement of our theorem, we introduce the notations $\mathcal{M}_{\text{comp},K}$ and $p_{\text{comp},K}$, which generalize $\mathcal{M}_{\text{comp}}$ and $p_{\text{comp}}$ that appear in Assumption 2.1. Let $\mathcal{M}_{\text{comp},K}$ denote the probability distribution of the comparison result obtained with $K$ independent LLM calls followed by majority voting (with ties broken randomly). Then we have
@@
Theorem 2.4.
Suppose that Assumption 2.1 holds and $N$ is a power of 2. Let $p_{i}$ be the probability that a candidate solution at the $i$-th level of the knockout tournament is correct, where $i=0,1,\dots,\log_{2}N$. [^1] In our notations, the zeroth level of the tournament contains $N$ initial candidates, the first level contains $N/2$ winners after the first round of pairwise comparisons, and so on. All candidates within the same level of the knockout tournament have the same probability of being a correct one, due to their symmetric roles. Then $p_{0}=p_{\text{gen}}$ and
@@
Consequently, the success probability of the overall algorithm, namely $p_{\log_{2}N}$, converges to 1 as $N$ grows; for any $0<\delta<0.5$, one has $p_{\log_{2}N}\geq 1-\delta$ as long as
@@
A proof can be found in Appendix B.2. The linear relationship between $\log_{2}N$ and $\log(1/\delta)$ reveals a power-law relationship between the failure probability $\delta$ and the number of candidates $N$.
@@ 3 A two-stage league-style algorithm
Algorithm 1 The proposed league-style algorithm
Input: the problem $x$.
 1. Generate NN candidates $y_{1},\dots,y_{N}\sim\mathcal{M}_{\text{gen}}(x)$.
 2. Compare each candidate $y_{i}$ against $K$ random opponents and estimate its average win rate $\widehat{\mu}_{i}$ by Eq. (2).
Output: the candidate with index $\widehat{i}\coloneqq\arg\max_{i}\widehat{\mu}_{i}$.
In this section, we propose a two-stage league-style algorithm that also enjoys a provable inference scaling law, under technical assumptions that are analogous to but more robust than those required by the knockout-style algorithm.
@@
[…]
$\widehat{\mu}_{i}\coloneqq\frac{1}{K}\sum_{j\in[K]}\phi\big(r_{i}(j),y_{i},y_{o_{i}(j)}\big),$ (2)
where $\phi(r_{i}(j),y_{i},y_{o_{i}(j)})$ denotes the score assigned, based on $r_{i}(j)$, to the candidate $y_{i}$ in its comparison against $y_{o_{i}(j)}$, e.g., 1 for a win, 0 for a loss, and 0.5 for a tie. Finally, the candidate with the highest average win rate $\widehat{\mu}_{i}$ is chosen (with ties broken randomly) as the output of the algorithm. See Algorithm 1 for a summary of this method.
@@
In other words, we assume that the LLM can generate, with non-zero probability, a correct solution whose average win rate against $\mathcal{M}_{\text{gen}}$ is higher than that of any incorrect solution; such a solution is called correct-and-strong by our definition. We also allow the existence of correct-but-weak solutions, imposing no assumption on their average win rates. Note that Assumption 3.1 can be tolerant of systematic errors by LLMs in comparing certain pairs of candidates, i.e., it may still hold true when there exist a correct solution $y$ and incorrect solution $y^{\prime}$ such that ${\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime})}\phi(r,y,y^{\prime})<0.5$, whereas Assumption 2.1 fails in such cases.
Remark 3.2.
One limitation of Assumption 3.1 is that it can be broken by an adversarial incorrect solution whose average win rate is unusually high, similar to the failure mode of best-of-N sampling discussed in Section 1. Nonetheless, on the presumption (backed by common practice and extensive empirical evidence) that pairwise comparison is more accurate and reliable than individual point-wise verification, we might safely say that Assumption 3.1 is conceptually weaker and more robust than the condition required by BoN (e.g., a perfect point-wise verifier) for provable inference scaling laws.
Intuitively, if Assumption 3.1 holds true and the hyperparameters $N$ and $K$ are sufficiently large, then with high probability, (1) there exist initial candidates that are correct-and-strong solutions, and (2) $\widehat{\mu}_{i}$ is an accurate estimate of $\mu_{y_{i}}$ for each $i\in[N]$. These conditions together lead to a correct final output of the algorithm. We formalize this intuition in the following theorem.
Theorem 3.3.
If Assumption 3.1 holds for the input problem, then the probability that the league-style algorithm (with hyperparameters $N$ and $K$) returns an incorrect final output is bounded by
@@
This theorem, whose proof can be found in Appendix B.3, ensures that the failure probability of the league-style algorithm decays to zero exponentially with respect to $N$ and $K$. Another way to interpret this theorem is as follows: to guarantee success probability $1-\delta$, it suffices to have
@@
Remark 3.4.
The provable success of the knockout-style algorithm relies on Assumption 2.1, while that of the league-style algorithm relies on Assumption 3.1. Although the latter is conceptually more robust than the former, we note that neither assumption is strictly weaker than the other (and thus both algorithms have their unique values). In other words, there exist scenarios where Assumption 2.1 holds true while Assumption 3.1 does not, and also scenarios where the reverse is true. Interested readers may refer to Appendix C for some minimal examples.
@@ 4 Experiments
We use three datasets for our experiments: GPQA rein2024gpqa , MMLU-Pro wang2024mmlu and MATH-500 math500 . GPQA consists of over 1000 graduate-level multiple-choice questions splitted into three categories (“main”, “diamond” and “extended”), all of which are used in our experiments. MMLU-Pro contains 14 categories of multiple-choice questions, some of which require advanced reasoning while others are more knowledge-focused. Due to limited computational resources, we use a randomly sampled subset of 100 questions for each category of MMLU-Pro in our experiments, which leads to a total of 1400 questions; we refer to this subset as MMLU-Pro-S throughout this work. MATH-500 is a subset of 500 problems from the MATH dataset introduced in lightman2024lets . Due to space limitations, we focus mainly on GPQA in this section, deferring empirical results for MMLU-Pro-S and MATH-500 (as well as supplementary results for GPQA) to Appendix D.
@@
We use Llama3.1-70B-Instruct (Llama3.1 for short) meta2024llama3herdmodels and Qwen2.5-72B-Instruct (Qwen2.5 for short) qwen2024qwen25technicalreport in our experiments, as well as a Mixed option that uses a mixture of both LLMs wang2024mixture ; zhang2024scalingllminferenceoptimized ; jiang2023llmblender : during the generation stage, half of the initial candidates are sampled by Llama3.1 and the other half by Qwen2.5; similarly, when a pair of candidates are compared for multiple times during the aggregation stage, half of them are done by Llama3.1 and the other half by Qwen2.5. The rationale is that the capabilities of different LLMs can be complementary to some extent, and thus using a mixture of them can make it more likely that Assumptions 2.1 and 3.1 hold true [^2] To formalize this intuition, consider a minimal scenario with two LLMs denoted by $M_{1}$ and $M_{2}$, and two problems denoted by $x_{1}$ and $x_{2}$. Suppose that $M_{1}$ is effective for the first problem $x_{1}$ (with $p_{\text{gen}}=0.2$ and $p_{\text{comp}}=0.7$) but ineffective for $x_{2}$ (with $p_{\text{gen}}=0$ and $p_{\text{comp}}=0.5$), while the reverse holds true for $M_{2}$. When either LLM is used alone, only one problem satisfies Assumption 2.1. However, when a mixture of two LLMs is used, both problems now satisfy Assumption 2.1 with $p_{\text{gen}}=(0+0.2)/2=0.1>0$ and $p_{\text{comp}}=(0.5+0.7)/2=0.6>0.5$, and thus can be solved by our algorithm. . Other models considered in our experiments include QwQ-32B qwq32b , a long-CoT reasoning LLM, and GPT-4o gpt4o , a proprietary API-based LLM; due to high computational or monetary costs, they are tested only on GPQA-diamond for a smaller range of $N$.
We leverage zero-shot chain-of-thought prompting Kojima2022ZeroShot for both generation and aggregation stages of the proposed algorithms. Unless specified otherwise, for the knockout-style algorithm, we fix $K=4$ for Llama3.1/ Qwen2.5/ Mixed, and $K=2$ for GPT-4o/ QwQ-32B; for the league-style algorithm, we consider a round-robin roundrobin version of it, with $K=4$ comparisons conducted between each of $N\choose 2$} pairs of initial candidates. To make the proposed algorithms efficient and scalable in practice, we implement them based on AgentScope gao2024agentscope , a multi-agent framework that supports parallel and distributed computation [^3] Our implementations can be found at https://github.com/pan-x-c/AgentScope/tree/feature/pxc/paper_provable/examples/paper_provable_scaling_law. Further implementation details can be found in Appendix D.1.